f(x)=3X+2 and g(x)=-2X+1 find g(f(-2))
I don't understand how to solve this problem.

Answer by ikleyn(52797)

.

First calculate f(-2) = 3*(-2) + 2 = -6 + 2 = -4.


After that, calculate  g(f(-2)),  replacing  f(-2)  here by (-4).


You will get  g(f(-2)) = g(-4) = -2*(-4) + 1 = 8+1 = 9.    ANSWER

Solved.

At this point, calculations are completed and the problem is solved, in full.


Never say again that you can not solve it.